The condensed income statement for the Blossom and Paul partnership for 2020 is as follows. Sales (240,000 units) $1,200,000 Cost of goods sold 800,000 Gross profit 400,000 Operating expenses Selling $280,000 Administrative 156,000 436,000 Net loss $(36,000 )

A cost behavior analysis indicates that 75% of the cost of goods sold are variable, 42% of the selling expenses are variable, and 40% of the administrative expenses are variable.

1) Compute the break-even point in total sales dollars for 2020.

2) Blossom has proposed a plan to get the partnership “out of the red” and improve its profitability. She feels that the quality of the product could be substantially improved by spending $0.30 more per unit on better raw materials. The selling price per unit could be increased to only $5.25 because of competitive pressures. Blossom estimates that sales volume will increase by 25%. Compute the net income under Blossom's proposal and the break-even point in dollars
Answer #1

Answer-1:

Variable cost: COGS 600,000 Selling expense Administrative expense 117,600 62,400 Total 780,000 %24 %24

Fixed cost: $ 200,000 COGS Selling expense Administrative expense 162,400 93,600 $ 456,000 Total

$ 1,200,000 Sales Less: Variable cost 780,000 Contribution margin 420,000

Contribution margin ratio = Contribution margin / Sales × 100

= 420,000 / 1,200,000 × 100 = 35%

Break-even point (in dollar) = Fixed cost / Contribution margin ratio

= 456,000 / 35% = $1,302,857

Answer-2:

Blossoms Plan $ 1,575,000 Sales Less: Variable cost: Cost of goods sold 840,000 Selling expense Administrative expense 147,0

Contribution margin ratio = Contribution margin / Sales × 100

= $510,000 / 1,575,000 × 100 = 32.38%

Break-even point (in dollar) = Fixed cost / Contribution margin ratio

= 456,000 / 32.38% = $1,408,277